OpenStudy (anonymous):

You have a coupon for your favorite clothing store for $25 off any purchase of more than $50. The store is also running a 20%-off sale on its entire inventory. Let x be the original price, be the price with the $25 coupon applied, and be the price with the 20% discount applied. A) Write an expression for f(x) B) Write an expression for g(x) C) What would an expression for (f o g)(x) represent D) What would an expression for (g o f)(x) represent E).If the store allows you to apply both the 20% discount and the $25-off coupon, does it matter which you apply first? How do you know?

OpenStudy (shamil98):

For A and B, what do you think the expression would be? x is the original price f(x) is the price with the 20$ coupon applied.. so ...

OpenStudy (shamil98):

f(x) only applies if it is greater than or equal to 50 as well, so keep that in mind.

OpenStudy (shamil98):

25$ coupon *

OpenStudy (anonymous):

x+25>= 50?

OpenStudy (shamil98):

A coupon means discount. and no.

OpenStudy (shamil98):

f(x) = x - 25, if x >= 50

OpenStudy (shamil98):

Do you understand why?

OpenStudy (anonymous):

Oh, so when you spend $50 or more you get a $25 discount

OpenStudy (shamil98):

Yes.

OpenStudy (shamil98):

good good , now try the expression for g(x) what is 20% off in decimal terms?

OpenStudy (anonymous):

.2

OpenStudy (shamil98):

20% off = 0.80 of the original price so g(x) would be then.

OpenStudy (anonymous):

g(x)= .80x?

OpenStudy (shamil98):

yes!

OpenStudy (shamil98):

in terms of words what would (f o g) represent, you're applying g first and then f.

OpenStudy (anonymous):

So I could say, (f o g)(x) represents the total price after using 20% off? I have always been horrible for the (f o g) and (g o f)

OpenStudy (shamil98):

you're applying 20% off and then the coupon

OpenStudy (anonymous):

So that is what I would write as my answer?

OpenStudy (shamil98):

yes.

OpenStudy (shamil98):

using analysis skills what then would be (g o f) ?

OpenStudy (anonymous):

When you apply the coupon before you get the 20% discount

OpenStudy (shamil98):

yes